Adult tickets for a play cost $15 and child tickets cost $6. If there were 33 people at a performance and the theater collected $360 from ticket sales, how many children attended the play?
Answer by ptaylor(2052)

Let x=number of children attending
Then 33-x=number of adults, right?
Amount collected from children = $6x
Amount collected from adults=$15(33-x)
And we are told that the above amounts = $360------So:
$6x+$15(33-x)=$360 get rid of parens (distributive law)
$6x+$495-$15x=$360 subtract $495 from both sides
$6x+$495-$495-$15x=$360-$495 collect like terms
-$9x=-$135 divide both sides by -9
x=15-----------------------------------number of children attending
33-x=33-15=18----------------------------number of adults attending
CK
18+15=33
33=33
also
$6*15+$15*18=$360
$90+$270=$360
$360=$360

Hope this helps---ptaylor